OpenStudy (ny,ny):

write the following as a single logarithm 6logb x + 1/2logb y.

OpenStudy (ny,ny):

\[6\log_{b}x+\frac{ 1 }{ 2 } \log_{b} y \]

Nnesha (nnesha):

power rule: apply this one first number or variable which is multiplying by log become exponent of y \[\huge\rm x \log y = \log y^x\]

OpenStudy (ny,ny):

hm

OpenStudy (ny,ny):

log x ^6?

Nnesha (nnesha):